Session: Jun 2013 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Noto (L. Nicotra at Fri May 24 07:07:06 2013) Success:

Torun (Kaz Borkowski at Fri May 24 07:34:25 2013) Success:
No known problems.

Medicina (C.Bortolotti at Fri May 24 08:12:44 2013) Success:
No know problems

Yebes (Daniel Valero at Fri May 24 08:23:51 2013) Success:
Not known problems

Shanghai (Bo Xia at Fri May 24 08:31:24 2013) Success:

Svetloe (Isaenko at Fri May 24 10:47:15 2013) Success:

Zelenchukskaya (Lysenkova M. at Fri May 24 10:47:36 2013) Success:

Badary (Zadrutsky, Sorokovikov at Fri May 24 10:49:46 2013) Minor failures: RECO
Scans No0036-No0044,No0054-No0062,No0108-No0116 failed due to Mark5 disk pack problems.

Urumqi (Peng Li at Fri May 24 11:55:06 2013) Success:
About the last one hour is not observed, because other tasks

Onsala (Michael Lindqvist at Sun May 26 13:17:00 2013) Success:
No known problems.

Effelsberg (Georgi at Mon May 27 08:02:02 2013) Success:

Westerbork (Geert Kuper at Mon May 27 19:10:29 2013) Success:
Wb had a problem with local correlated data storage due to a broken local disk. Therefore we have only part of the cal file (missing from start of project till 06:55 UT .See comments in cal file) . The VLBI data should be ok, altough we could not check it.